# MC-MoE: Mixture Compressor for Mixture-of-Experts LLMs Gains More

### v1

@@ Abstract
Mixture-of-Experts large language models (MoE-LLMs) marks a significant step forward of language models, however, they encounter two critical challenges in practice: 1) expert parameters lead to considerable memory consumption and loading latency; and 2) the current activated experts are redundant, as many tokens may only require a single expert. Motivated by these issues, we investigate the MoE-LLMs and make two key observations: a) different experts exhibit varying behaviors on activation reconstruction error, routing scores, and activated frequencies, highlighting their differing importance, and b) not all tokens are equally important– only a small subset is critical. Building on these insights, we propose MC-MoE, a training-free Mixture-Compressor for MoE-LLMs, which leverages the significance of both experts and tokens to achieve an extreme compression. First, to mitigate storage and loading overheads, we introduce Pre-Loading Mixed-Precision Quantization (PMQ), which formulates the adaptive bit-width allocation as a Linear Programming (LP) problem, where the objective function balances multi-factors reflecting the importance of each expert. Additionally, we develop Online Dynamic Pruning (ODP), which identifies important tokens to retain and dynamically select activated experts for other tokens during inference to optimize efficiency while maintaining performance. Our MC-MoE integrates static quantization and dynamic pruning to collaboratively achieve extreme compression for MoE-LLMs with less accuracy loss, ensuring an optimal trade-off between performance and efficiency. Extensive experiments confirm the effectiveness of our approach. For instance, at 2.54 bits, MC-MoE compresses 76.6% of the model, with only a 3.8% average accuracy loss. During dynamic inference, we further reduce activated parameters by 15%, with a performance drop of less than 0.6%. Remarkably, MC-MoE even surpasses floating-point 13b dense LLMs with significantly smaller parameter sizes, suggesting that mixture compression in MoE-LLMs has the potential to outperform both comparable and larger dense LLMs. Our code is available at https://github.com/Aaronhuang-778/MC-MoE.
@@ 1 Introduction
To this end, we propose the MC-MoE, i.e., Mixture-Compressor for MoE LLMs, exploring the combined benefits of expert quantization and pruning. MC-MoE consists of two phases: Pre-Loading Mixed-Precision Quantization (PMQ) and Online Dynamic Pruning (ODP), as shown in Fig. 1(a). In the pre-loading phase, we focus on extreme compression of the stored experts through low-bit quantization. Our empirical study reveals imbalances in activation reconstruction error, routing weights, and frequencies of activated expert (Sec. 3.2.1 and Fig. 3), which inspires the allocation of different bit-widths to each expert. However, relying solely on the routing frequencies or scores is insufficient to accurately determine the optimal bit-width, as the two distributions may not be consistent but rather the opposite (Li et al., 2024). Therefore, we developed a weighted evaluation function that considers both the frequency and scores of expert activations, as well as the associated quantization loss at different bit-widths. This function is then minimized within a Linear Programming (LP) model to determine the optimal quantization configuration. Utilizing a training-free Post-training Quantization (PTQ) approach, GPTQ (Frantar et al., 2022), PMQ achieves high-performance compression at extremely low bit-widths (1.5$\sim$2.5-bit), and our mixed-precision strategy is compatible with other advanced quantization techniques (Tseng et al., 2024; Chen et al., 2024; Shao et al., 2023; Egiazarian et al., 2024; Liao & Monz, 2024). As for inference phase, ODP dynamically prunes low-confidence experts for each token based on the routing weights. Our pruning strategy follows two key principles: first, experts with significantly lower routing scores are categorized as ”low confidence” and can be pruned (Lu et al., 2024). Second, to prevent attention degradation that solely relies on routing weights, we protect important tokens by considering both attention scores and feature magnitudes. Experiments show that protecting only 2% of the important tokens effectively mitigates pruning loss while maintaining nearly the same compression ratio.
@@ 2 Related Works
Mixture-of-Experts LLMs. LLMs have achieved significant advancements across various natural language processing domains (Chang et al., 2024; Zhao et al., 2023). Despite their success, these models rely heavily on dense parameters, which presents significant challenges for deployment(Zhou et al., 2024; Zhu et al., 2023). Sparse activated MoE models have been identified as an essential strategy to enhance the cost-performance balance in LLMs. In MoE models, each layer is comprised of several experts, with each token activating only a specific subset, thereby significantly improving efficiency compared to dense models, which activate all parameters for every input (Shazeer et al., 2017; Yun et al., 2024). Recent advancements in decoder-only LLMs (Brown, 2020) have further popularized MoE-based architectures (Jiang et al., 2024; Muennighoff et al., 2024). Industry-leading models such as Mixtral $8\times 7$(Jiang et al., 2024), Deepseek-MoE (Dai et al., 2024), and GPT-4 (Achiam et al., 2023) also incorporate this technology.
Quantization for LLMs. Post-Training Quantization (PTQ) is an efficient method that requires no additional training, making it well-suited for large-scale LLMs (Dettmers et al., 2022; Frantar et al., 2022; Xiao et al., 2023; Shao et al., 2023; Lin et al., 2024). Previous studies have investigated the diverse salience of weights and proposed mixed-precision methods to improve low-bitwidth performance by allocating different bitwidths accordingly (Dong et al., 2020; Huang et al., 2024c; Dettmers et al., 2023; Shang et al., 2023; Huang et al., 2024a). Recent research introduced an expert-guided, block-wise mixed-precision benchmark for MoE-LLMs to address the disparities in expert weights (Li et al., 2024); however, developing more effective expert-wise quantization strategies remains a challenge. Codebook-based encoding approaches enable more precise quantization of LLMs and enhance post-quantization performance through fine-tuning (Egiazarian et al., 2024; Tseng et al., 2024). While Quantization Aware Training (QAT) requires significant resources (Chen et al., 2024; Liu et al., 2023), QAT-based retraining strategies or PTQ combined with additional fine-tuning (Liao & Monz, 2024; Guo et al., 2023; Huang et al., 2024b) are more effective in maintaining the performance of quantized lightweight LLMs.
@@
Figure 2: The overview of our proposed MC-MoE pipeline with two stages compression for experts. (a) Framework of pre-loading static mixed-precision quantization (PMQ) of MoE-LLMs. PMQ determins the activated feature and loss sensitivity of all experts and plans the optimal precision configuration under ultra-low -bit-width. (b) Schematic of online dynamic mixture pruning (ODP) of MoE-LLMs. ODP utilizes significant token protection mechanism with weigh-guided experts pruning, which only need to keep 2% token to successfully safeguard the MoE performance.
@@ 3.1 Preliminaries
where $\mathbf{W}^{q}$ denotes the quantized weight and $\|\cdot\|_{2}$ is the mean square error (MSE) loss. The primary objective of this study is to explore the optimal mixture compression strategy for MoE-LLMs. To this end, we employ the efficient PTQ scheme, GPTQ (Frantar et al., 2022), as our foundational tool. By utilizing Hessian-based estimation ($\mathbf{H}=2\mathbf{X}\mathbf{X}^{\top}$) and quantization error compensation, GPTQ effectively reduces the group-wise quantization error of weights, enabling the quantization of Mixtral $8\times 7$b within 90 minutes. This work focuses on the design of optimal mixture compression strategies for MoE-LLMs and is therefore orthogonal to other quantization techniques, including PTQ methods (Shao et al., 2023; Lin et al., 2024), codebook-based works (Egiazarian et al., 2024; Tseng et al., 2024), and even the deployment of fine-tuning (Liao & Monz, 2024) or QAT (Chen et al., 2024; Liu et al., 2023) can be deployed for MC-MoE, additional evidences are shown in Appendix A.3.
@@ 3.2.1 Experts Significance Analysis
The core principle of our expert quantization strategy is grounded in the significance of each expert, which enables the allocation of bit-widths according to their relative importance within a block. We initially observed the performance of different experts in Mixtral $8\times 7$b in terms of expert-drop reconstruction loss (Frobenius norm) (He et al., 2017), and activation features on the dataset C4 (Raffel et al., 2020) and the specialized domain dataset Math (Hendrycks et al., 2021). As shown in Fig 3, the impact of experts on the model varies widely: 1) some experts, such as the one at position $[2,4]$, have minimal influence on the output activation reconstruction loss, while others in layer 2 exhibit significantly higher losses, highlighting the imbalance among MoE-LLMs’ experts; 2) the activation scores and frequencies reveal distinct patterns, where experts at positions $[11,3]$ and $[12,7]$ show extremely low activation frequencies and average scores, while the expert at position $[1,3]$ has low scores but comparatively high activation frequencies; and 3) in task-specific contexts like mathematics (MATH), MoE activates fewer experts, resulting in a strong sparser distribution than general text datasets. This variability in routing feature inspires the need to consider multiple factors in determining the optimal bit-width allocation for experts.
We mainly measure the significance of each expert through two critical factors: access frequency and activation weight. Given an $N$-sized calibration dataset C4 (general language understanding dataset), we first perform inference on the original 16-bit MoE-LLMs to compute these factors. For each expert, access frequency refers to the rate at which the expert is activated, defined as the number of activations divided by the dataset size. Thus, $i$-th expert’s access frequency is $\phi_{i}=\frac{n_{i}}{N}$, where $n_{i}$ is this expert’s total activated number. A higher activation frequency indicates that the expert is triggered more often, suggesting its generality and applicability across a wide range of tokens, thereby signifying higher importance. However, access frequency alone overlooks the potential significance of experts who are rarely activated. To account for this, we introduce the activation-weighted metric, which sums the routing weights assigned to each expert during inference. This metric for $i$-th expert can be denoted as $w_{i}=\frac{\sum_{j=1}^{N}{\sigma_{j}}}{N}$, where $\sigma_{j}$ is the expert’s routing weight in the $j$-th inference. This provides a finer-grained measure of an expert’s contribution in MoE-LLMs, capturing its relative importance beyond mere frequencies. By combining these two metrics, we achieve a more nuanced assessment of expert significance. The final expert significance is computed as $\phi_{i}^{\alpha}\cdot w_{i}^{\beta}$, where $\alpha$ and $\beta$ are hyperparameters used to balance the two factors.
@@ 3.3 Online Dynamic Pruning
Our PMQ strategy compresses the storage memory of experts during the pre-loading phase; however, the selection of the $top\mbox{-}k$ experts during inference still incurs high computational costs. As discussed in Lu et al. (2024), not all tokens require $k$ experts for inference. Moreover, quantization, as a static compression strategy, struggles to transform the precision of experts during inference. To optimize efficiency while maintaining performance during inference, in this section, we introduce the Online Dynamic Pruning (ODP) technique, which identifies important tokens to retain and dynamically selects activated experts for other tokens.
@@ 3.3.1 Attention Decay Under Weight-only Pruning
$w_{0}$ and $w_{1}$ denote the $top\mbox{-}2$ experts, respectively, with $\mu$ erving as a hyperparameter threshold for each MoE layer. This threshold is set at the median value of $\frac{w_{1}}{w_{0}}$ derived from calibration data (Lu et al., 2024). According to Eq. (5), when a selected expert has a notably low weight, it is feasible to be pruned for the current token, thus retaining only the primary expert for computation. Sec. 4.2 documents that employing this weight-based dynamic pruning strategy reduces computational demands by 15%, but also incurs a performance decrease of approximately 10%. Further examination reveals that this decline is due to an ”attention decay” effect, which is evident in Fig. 4. Specifically, unpruned conditions show a pronounced vertical pattern in the attention map of block 15, head 4, at token 67 (Fig. 4, middle). However, the application of weight-only pruning in block 14 results in the elimination of one expert for token 67, leading to a significant reduction in the attention map score at token 67 in block 15 (Fig. 4, left). This effect is herein defined as ”attention decay.”
@@ 3.3.2 Significance-Aware Token Protection
As demonstrated in Fig. 4, right, introducing important token protection into weight-only pruning effectively mitigates attention decay issues. Due to the high importance parameter $I_{67}$ of token 67, all experts are preserved for computing token 67 in block 14, thereby preserving the expected distribution in the attention map of block 15 for token 67. Experiments in Sec. 4.2 indicate that selectively protecting merely 2% of important tokens can significantly reduce performance losses in MoE-LLMs, while still maintaining a computational efficiency improvement of approximately 15%.
@@ 4 Experiment
In this section, a series of experiments are conducted to evaluate the proposed MC-MoE. We present by describing the models selected, datasets, quantization methods, and experimental parameter settings. In Sec. 4.1, we assess the parameter settings for the PMQ method and the performance of mixture quantization. We conduct a detailed evaluation of the performance loss and compression efficiency of ODP stage, shown in Sec. 4.2. Finally, we present the combined performance of static quantization and dynamic pruning in MoE-LLMs.
@@ 4.1 Experiment on Pre-Loading Mixed-Precision Quantization
Table 2: Performance of quantized Mixtral $8\times 7$b on eight zero-shot benchmarks. We deploy GPTQ as our baseline PTQ method for uniform quantization. ”Uni” denotes the uniform quantization of 2-bit with GPTQ. Since the results of some data sets in the block score predictor (BSP) (Li et al., 2024) method were not reported, we resumed the relevant quantized model from the official code repository and evaluated all the results under the same settings. In BSP, 25% MoE layers are 4-bit and the left are 2-bit to achieve 2.54-bit. ”HellaS.” is the short format of ”HellaSwag” and ”Wino.” denotes ”Winogrande”. $\downarrow$ gives the accuracy loss between quantized results and original 16-bit model.
@@
Figure 7: Significant tokens protection of 2.05-bit Mixtral $8\times 7$b. “CR” denotes the average computation compression ratio (blue); ”PPL” denotes the perplexity(red); Star represents the weight-only pruning performance.
Figure 8: Less significant tokens drop of 2.05-bit Mixtral $8\times 7$b. In Mixtral $8\times 7$b, we mask all experts of the less significant tokens. “CR” denotes the average computation compression ratio (blue); ”PPL” denotes the perplexity(red).
Comparison of Mixed-Precision Quantization. We present a comprehensive comparison of the performance of PMQ within the ultra-low bit-width range. GPTQ was set as the baseline for uniform bit-width quantization, denoted as ”Uni” in Tab. 2. We also compare it with a recent mixed-precision approach for MoE-LLMs known as the block score predictor (BSP) (Li et al., 2024). Following Eq. 4, we set the average bit-width of Mixtral $8\times 7$b within the range of 1.57-bit to 2.54-bit. As shown in Tab. 2, the uncompressed 16-bit model achieves an average accuracy of 71.29%. With uniform precision quantization, the average loss for the 3-bit model is approximately 2.2%, while the loss for the 2-bit model increases significantly by 28.6%, highlighting the challenges in maintaining model accuracy with existing uniform precision quantization methods at ultra-low bit widths.
@@
Table 4: Ablation evaluation of PMQ and ODP for Mixtral $8\times 7/22$b, and compared with dense LLaMA models. ”Params” denotes the parameter size, and ”Act Params” is averaged activated parameters for one token. The parameter calculation of the compressed model includes the compressed weights and quantizer parameters (e.g., scaling factor and zero factor for dequantization). We carry out the average activated parameter size and speedup on C4 dataset. 16-bit Mixtral $8\times 7$b uses 2 A100-80GB GPUs and Mixtral $8\times 22$b uses 4, quantized models are tested on one A100-80GB GPU.

### v2

# Mixture Compressor for Mixture-of-Experts LLMs Gains More

@@ Abstract
Mixture-of-Experts large language models (MoE-LLMs) marks a significant step forward of language models, however, they encounter two critical challenges in practice: 1) expert parameters lead to considerable memory consumption and loading latency; and 2) the current activated experts are redundant, as many tokens may only require a single expert. Motivated by these issues, we investigate the MoE-LLMs and make two key observations: a) different experts exhibit varying behaviors on activation reconstruction error, routing scores, and activated frequencies, highlighting their differing importance, and b) not all tokens are equally important– only a small subset is critical. Building on these insights, we propose MC, a training-free Mixture-Compressor for MoE-LLMs, which leverages the significance of both experts and tokens to achieve an extreme compression. First, to mitigate storage and loading overheads, we introduce Pre-Loading Mixed-Precision Quantization (PMQ), which formulates the adaptive bit-width allocation as a Linear Programming (LP) problem, where the objective function balances multi-factors reflecting the importance of each expert. Additionally, we develop Online Dynamic Pruning (ODP), which identifies important tokens to retain and dynamically select activated experts for other tokens during inference to optimize efficiency while maintaining performance. Our MC integrates static quantization and dynamic pruning to collaboratively achieve extreme compression for MoE-LLMs with less accuracy loss, ensuring an optimal trade-off between performance and efficiency. Extensive experiments confirm the effectiveness of our approach. For instance, at 2.54 bits, MC compresses 76.6% of the model, with only a 3.8% average accuracy loss in eight commonsense benchmarks. During dynamic inference, we further reduce activated parameters by 15%, with a performance drop of less than 0.6%. Remarkably, MC even surpasses floating-point 13b dense LLMs with significantly smaller parameter sizes, suggesting that mixture compression in MoE-LLMs has the potential to outperform both comparable and larger dense LLMs. Our code is available at https://github.com/Aaronhuang-778/MC-MoE.
@@ 1 Introduction
To this end, we propose the MC, i.e., Mixture-Compressor for MoE LLMs, exploring the combined benefits of expert quantization and pruning. MC consists of two phases: Pre-Loading Mixed-Precision Quantization (PMQ) and Online Dynamic Pruning (ODP), as shown in Fig. 1(a). In the pre-loading phase, we focus on extreme compression of the stored experts through low-bit quantization. Our empirical study reveals imbalances in activation reconstruction error, routing weights, and frequencies of activated expert (Sec. 3.2.1 and Fig. 3), which inspires the allocation of different bit-widths to each expert. However, relying solely on the routing frequencies or scores is insufficient to accurately determine the optimal bit-width, as the two distributions may not be consistent but rather the opposite (Li et al., 2024). Therefore, we developed a weighted evaluation function that considers both the frequency and scores of expert activations, as well as the associated quantization loss at different bit-widths. This function is then minimized within a Linear Programming (LP) model to determine the optimal quantization configuration. Utilizing a training-free Post-training Quantization (PTQ) approach, GPTQ (Frantar et al., 2022), PMQ achieves high-performance compression at extremely low bit-widths (1.5$\sim$2.5-bit), and our mixed-precision strategy is compatible with other advanced quantization techniques (Tseng et al., 2024; Chen et al., 2024; Shao et al., 2023; Egiazarian et al., 2024; Liao & Monz, 2024). As for inference phase, ODP dynamically prunes low-confidence experts for each token based on the routing weights. Our pruning strategy follows two key principles: first, experts with significantly lower routing scores are categorized as “low confidence” and can be pruned (Lu et al., 2024). Second, to prevent attention degradation that solely relies on routing weights, we protect important tokens by considering both attention scores and feature magnitudes. Experiments show that protecting only 2% of the important tokens effectively mitigates pruning loss while maintaining nearly the same compression ratio.
@@ 2 Related Works
Mixture-of-Experts LLMs. LLMs have achieved significant advancements across various natural language domains (Chang et al., 2024; Zhao et al., 2023). Despite their success, these models rely heavily on dense parameters, which presents significant challenges for deployment(Zhou et al., 2024; Zhu et al., 2023). Sparse activated MoE models have been identified as an essential strategy to enhance the cost-performance balance in LLMs. In MoE models, each layer is comprised of several experts, with each token activating only a specific subset, thereby significantly improving efficiency compared to dense models, which activate all parameters for every input (Shazeer et al., 2017; Yun et al., 2024). Recent advancements in LLMs (Brown, 2020) have further popularized MoE-based architectures (Jiang et al., 2024; Muennighoff et al., 2024). Industry-leading models such as Mixtral $8\times 7$(Jiang et al., 2024) and Deepseek-R1 (Guo et al., 2025) also incorporate this technology.
Quantization for LLMs. Post-Training Quantization (PTQ) is an efficient method that requires no additional training, making it well-suited for large-scale LLMs (Dettmers et al., 2022; Frantar et al., 2022; Xiao et al., 2023; Shao et al., 2023; Lin et al., 2024). Previous studies have investigated the diverse salience of weights and proposed mixed-precision methods to improve low-bitwidth performance by allocating different bitwidths accordingly (Dong et al., 2020; Huang et al., 2024c; Dettmers et al., 2023; Shang et al., 2023; Huang et al., 2024a). Recent research introduced an expert-guided, block-wise mixed-precision benchmark for MoE-LLMs to address the disparities in expert weights (Li et al., 2024); however, developing more effective expert-wise quantization strategies remains a challenge. Codebook-based encoding approaches enable more precise quantization of LLMs and enhance post-quantization performance through fine-tuning (Egiazarian et al., 2024; Tseng et al., 2024). While Quantization Aware Training (QAT) requires significant resources (Chen et al., 2024; Liu et al., 2023b), QAT-based retraining strategies or PTQ combined with additional fine-tuning (Liao & Monz, 2024; Guo et al., 2023; Huang et al., 2024b) are more effective in maintaining the performance of quantized lightweight LLMs.
@@
Figure 2: The overview of our proposed MC pipeline with two stages compression for experts. (a) Framework of pre-loading static mixed-precision quantization (PMQ) of MoE-LLMs. PMQ determins the activated feature and loss sensitivity of all experts and plans the optimal precision configuration under ultra-low -bit-width. (b) Schematic of online dynamic mixture pruning (ODP) of MoE-LLMs. ODP utilizes significant token protection mechanism with weigh-guided experts pruning, which only need to keep 2% token to successfully safeguard the MoE performance.
@@ 3.1 Preliminaries
where $\mathbf{W}^{q}$ denotes the quantized weight and $\|\cdot\|_{2}$ is the mean square error (MSE) loss. The primary objective of this study is to explore the optimal mixture compression strategy for MoE-LLMs. To this end, we employ the efficient PTQ scheme, GPTQ (Frantar et al., 2022), as our foundational tool. By utilizing Hessian-based estimation ($\mathbf{H}=2\mathbf{X}\mathbf{X}^{\top}$) and quantization error compensation, GPTQ effectively reduces the group-wise quantization error of weights, enabling the quantization of Mixtral $8\times 7$b within 90 minutes. This work focuses on the design of optimal mixture compression strategies for MoE-LLMs and is therefore orthogonal to other quantization techniques, including PTQ methods (Shao et al., 2023; Lin et al., 2024), codebook-based works (Egiazarian et al., 2024; Tseng et al., 2024), and even the deployment of fine-tuning (Liao & Monz, 2024) or QAT (Chen et al., 2024; Liu et al., 2023b) can be deployed for MC, additional evidences are shown in Appendix A.3.
@@ 3.2.1 Experts Significance Analysis
The core principle of our expert quantization strategy is grounded in the significance of each expert, which enables the allocation of bit-widths according to their relative importance within a block. We initially observed the performance of different experts in Mixtral $8\times 7$b in terms of expert-drop reconstruction loss (Frobenius norm) (He et al., 2017), and activation features on the dataset C4 (Raffel et al., 2020) and the specialized domain dataset Math (Hendrycks et al., 2021). As shown in Fig 3, the impact of experts on the model varies widely: 1) some experts, such as the one at position $[2,4]$ (Fig 3 left), have minimal influence on the output activation reconstruction loss, while others in layer 2 exhibit significantly higher losses, highlighting the imbalance among MoE-LLMs’ experts; 2) the activation scores and frequencies reveal distinct patterns, where experts at positions $[11,3]$ and $[12,7]$ show extremely low activation frequencies and average scores, while the expert at position $[1,3]$ has low scores but comparatively high activation frequencies; and 3) in task-specific contexts like mathematics, MoE activates fewer experts, resulting in a sparser distribution than general tasks. This variability in routing feature inspires the need to consider multiple factors in determining the optimal bit-width allocation for experts.
We mainly measure the significance of each expert through two factors: access frequency and activation weight. Given an $N$-sized calibration dataset C4 (general language understanding dataset), we first perform inference on the original 16-bit MoE-LLMs. For each expert, access frequency refers to the rate at which the expert is activated. Thus, $i$-th expert’s access frequency is $\phi_{i}=\frac{n_{i}}{N}$, where $n_{i}$ is this expert’s total activated number. A higher activation frequency indicates that the expert is triggered more often, suggesting its generality and applicability across a wide range of tokens. However, access frequency alone overlooks the potential significance of experts who are rarely activated. To account for this, we introduce the activation-weighted metric, which sums the routing weights assigned to each expert during inference. This metric for $i$-th expert can be denoted as $w_{i}=\frac{\sum_{j=1}^{N}{\sigma_{j}}}{N}$, where $\sigma_{j}$ is the expert’s routing weight in the $j$-th inference. This provides a finer-grained measure of an expert’s contribution in MoE-LLMs, capturing its relative importance beyond mere frequencies. The final expert significance is computed as $\phi_{i}^{\alpha}\cdot w_{i}^{\beta}$, where $\alpha$ and $\beta$ are hyperparameters used to balance the two factors.
@@ 3.3 Online Dynamic Pruning
Our PMQ strategy compresses the storage memory of experts during the pre-loading phase; however, the selection of the $top\mbox{-}k$ experts during inference still incurs high computational costs. As discussed in Lu et al. (2024), not all tokens require $k$ experts for inference. To optimize efficiency while maintaining performance during inference, in this section, we introduce the Online Dynamic Pruning (ODP) technique, which identifies important tokens to retain and dynamically selects activated experts for other tokens.
@@ 3.3.1 Attention Decay Under Weight-only Pruning
$w_{0}$ and $w_{1}$ denote the $top\mbox{-}2$ experts, respectively, with $\mu$ erving as a hyperparameter threshold for each MoE layer. This threshold is set at the median value of $\frac{w_{1}}{w_{0}}$ derived from calibration data (Lu et al., 2024). According to Eq. (5), when a selected expert has a notably low weight, it is feasible to be pruned for the current token, thus retaining only the primary expert for computation. Sec. 4.2 documents that employing this weight-based dynamic pruning strategy reduces computational demands by 15%, but also incurs a performance decrease of approximately 10%. Further examination reveals that this decline is due to an “attention decay” effect, which is evident in Fig. 4. Specifically, unpruned conditions show a pronounced vertical pattern in the attention map of block 15, head 4, at token 67 (Fig. 4, middle). However, the application of weight-only pruning in block 14 results in the elimination of one expert for token 67, leading to a significant reduction in the attention map score at token 67 in block 15 (Fig. 4, left). This effect is herein defined as “attention decay.”
@@ 3.3.2 Significance-Aware Token Protection
As demonstrated in Fig. 4, right, introducing important token protection into weight-only pruning effectively mitigates attention decay issues. Due to the high importance parameter $I_{67}$ of token 67, all experts are preserved for computing token 67 in block 14, thereby preserving the expected distribution in the attention map of block 15 for token 67. Experiments in Sec. 4.2 indicate that selectively protecting merely 2% of important tokens can significantly reduce performance losses in MoE-LLMs, while still maintaining a computational efficiency improvement of approximately 15%. We also provide the detailed computation overhead analysis in Appendix A.9.
@@ 4 Experiment
In this section, a series of experiments are conducted to evaluate the proposed MC. We present by describing the experimental parameter settings and results. In Sec. 4.1, we assess the parameter settings for the PMQ method and the performance of mixture quantization. We conduct a detailed evaluation of the performance loss and compression efficiency of ODP stage, shown in Sec. 4.2. Finally, we present the combined performance of MoE mixture compressor.
@@ 4.1 Experiment on Pre-Loading Mixed-Precision Quantization
Table 2: Performance of quantized Mixtral $8\times 7$b on eight zero-shot benchmarks. We deploy GPTQ as our baseline PTQ method for uniform quantization. “Uni” denotes the uniform quantization of 2-bit with GPTQ. Since the results of some data sets in the block score predictor (BSP) (Li et al., 2024) method were not reported, we resumed the relevant quantized model from the official code repository and evaluated all the results under the same settings. In BSP, 25% MoE layers are 4-bit and the left are 2-bit to achieve 2.54-bit. “HellaS.” is the short format of “HellaSwag” and “Wino.” denotes “Winogrande”. $\downarrow$ gives the accuracy loss between quantized results and original 16-bit model.
@@
Figure 7: Significant tokens protection of 2.05-bit Mixtral $8\times 7$b. “CR” denotes the average computation compression ratio (blue); “PPL” denotes the perplexity(red); Star represents the weight-only pruning performance.
Figure 8: Less significant tokens drop of 2.05-bit Mixtral $8\times 7$b. In Mixtral $8\times 7$b, we mask all experts of the less significant tokens. “CR” denotes the average computation compression ratio (blue); “PPL” denotes the perplexity(red).
Comparison of Mixed-Precision Quantization. We present a comprehensive comparison of the performance of PMQ within the ultra-low bit-width range. GPTQ was set as the baseline for uniform bit-width quantization, denoted as “Uni” in Tab. 2. We also compare it with a recent mixed-precision approach for MoE-LLMs known as the block score predictor (BSP) (Li et al., 2024). Following Eq. 4, we set the average bit-width of Mixtral $8\times 7$b within the range of 1.57-bit to 2.54-bit. As shown in Tab. 2, the uncompressed 16-bit model achieves an average accuracy of 71.29%. With uniform precision quantization, the average loss for the 3-bit model is approximately 2.2%, while the loss for the 2-bit model increases significantly by 28.6%, highlighting the challenges in maintaining model accuracy with existing uniform precision quantization methods at ultra-low bit widths.
@@
Table 4: Ablation evaluation of PMQ and ODP for Mixtral $8\times 7/22$b, and compared with dense LLaMA models. “Params” denotes the parameter size, and “Act Params” is averaged activated parameters for one token. The parameter calculation of the compressed model includes the compressed weights and quantizer parameters (e.g., scaling factor and zero factor for dequantization). We carry out the average activated parameter size and speedup on C4 dataset. 16-bit Mixtral $8\times 7$b uses 2 A100-80GB GPUs and Mixtral $8\times 22$b uses 4, quantized models are tested on one A100-80GB GPU.